OpenStudy (anonymous):

At the peak of a beanbag animal fad, one sales representative sold 12,000 of the animals in one month. Each month after that, the rep sold about 20% fewer animals. (a) Write the equation that represents the number of beanbag animals the sales representative sold. (b) About how many beanbag animals did the rep sell in the 6th month after the peak? I am really confused on how to find this... Could someone help?

OpenStudy (the_fizicx99):

Hmm, well the principle is 12k and the rate is 20% I'm thinking, f(x) = 12000(1-0.20)^x

OpenStudy (the_fizicx99):

15,500? where did they get 15,500?

OpenStudy (the_fizicx99):

I'm not sure how to answer given those options, though, for A i'm pretty sure its f(x) = 12000(1-0.20)^x. For B, f(6) = 12000(0.8)^6 = f(6) = 12000(0.262144) = 3,145.728 rounded to 3,146. So f(6) = 3,146

OpenStudy (anonymous):

I am sorry, I went back to the page and I read the wrong part. You were right thank you :)

OpenStudy (the_fizicx99):

xDD lol yw